2|x+1|-3=0
+3
2|x+1|=3
X+1=3/2
How can I turn the fraction into a whole number to continue solving the problem?

Answer by Alan3354(69443)

2|x+1|-3=0
+3
2|x+1|=3
X+1=3/2
How can I turn the fraction into a whole number to continue solving the problem?
----------------
3/2 can't be a whole number.
====================================
X+1=3/2
Subtract 1
x = 1/2 (1st solution)
================
|x+1| = 3/2
-(x+1) = 3/2
x+1 = -3/2
x = -5/2 (2nd solution)